# LaTeXTrans: Structured LaTeX Translation with Multi-Agent Coordination

### v2

@@ Abstract
Despite the remarkable progress of modern machine translation (MT) systems on general-domain texts, translating structured LaTeX-formatted documents remains a significant challenge. These documents typically interleave natural language with domain-specific syntax, such as mathematical equations, tables, figures, and cross-references, all of which must be accurately preserved to maintain semantic integrity and compilability. In this paper, we introduce LaTeXTrans, a collaborative multi-agent system designed to address this challenge. LaTeXTrans ensures format preservation, structural fidelity, and terminology consistency through six specialized agents: 1) a Parser that decomposes LaTeX into translation-friendly units via placeholder substitution and syntax filtering; 2) a Translator, Validator, Summarizer, and Terminology Extractor that work collaboratively to ensure context-aware, self-correcting, and terminology-consistent translations; 3) a Generator that reconstructs the translated content into well-structured LaTeX documents. Experimental results demonstrate that LaTeXTrans can outperform mainstream MT systems in both translation accuracy and structural fidelity, offering an effective and practical solution for translating LaTeX-formatted documents. The code of LaTeXTrans is available at https://github.com/NiuTrans/LaTeXTrans.
@@
Ziming Zhu1††thanks: Authors contributed equally., Chenglong Wang1∗, Shunjie Xing1, Yifu Huo1, Fengning Tian2, Quan Du2, Di Yang1,2, Chunliang Zhang1,2, Tong Xiao1,2††thanks: Corresponding author. and Jingbo Zhu1,2 1 School of Computer Science and Engineering, Northeastern University, Shenyang, China 2 NiuTrans Research, Shenyang, China {zhuzm0721, clwang1119}@gmail.com, {xiaotong, zhujingbo}@mail.neu.edu.cn
@@ 1 Introduction
LaTeX is a widely adopted macro package system built on top of TeX, designed to facilitate the typesetting of complex and structured documents. It has become the de facto standard for scholarly publications across a wide range of scientific disciplines. According to recent statistics, nearly 98% of scientific papers are published in English, while only about 3% of the global population speaks English as their first language (Kleidermacher and Zou, 2025). This linguistic disparity places considerable pressure on non-native English speakers, who are frequently required to read or write LaTeX-formatted documents in English. As a result, the technical barriers to academic learning and research are significantly increased.
A straightforward approach to ease this burden is to translate LaTeX documents into the user’s native language by processing the compiled PDF version, a process referred to as PDF translation. However, this approach often results in incomplete formatting due to errors in PDF parsing. In contrast, a more promising alternative is to translate directly at the LaTeX source level and then compile the translated content into a target-language PDF document. This approach can preserve structural information and allows better control over formatting.
However, translating LaTeX source files presents unique challenges not encountered in plain-text translation. LaTeX documents interleave natural language with domain-specific markup, such as mathematical equations, citation commands, and formatting environments, all of which must be precisely preserved to ensure semantic correctness and successful compilation. Naively applying standard MT systems to LaTeX code typically leads to broken syntax, semantic errors, or formatting loss, ultimately hindering rather than helping the user.
To address these challenges, in this paper, we introduce LaTeXTrans, a collaborative multi-agent system designed to directly translate LaTeX source files while preserving their structural and semantic integrity. Our LaTeXTrans operates on raw LaTeX code and maintains the full syntactic and semantic structure of the document throughout the entire translation pipeline. Specifically, it comprises three modules and six specialized agents:
Parsing Module: Responsible for fine-grained analysis of LaTeX-formatted documents. To handle the structural complexity of LaTeX, we design a Parser agent equipped with a placeholder mechanism and a syntax filter, which together decompose the source into manageable translation units.
@@
To evaluate the effectiveness of LaTeXTrans, we first construct a LaTeX source test set using TeX files collected from arXiv papers. We then compare LaTeXTrans with a range of MT and LLM-based translation baselines. Experimental results demonstrate that LaTeXTrans consistently outperforms all baselines in both translation accuracy and format fidelity. Notably, LaTeXTrans achieves an improvement of 13.20 points on FC-score, along with significant gains in COMETkiwi and LLM-score when compared to GPT-4o.
@@ 2 Related works
The emergence of LLMs has introduced a new paradigm for MT, shifting away from traditional supervised learning on parallel corpora toward more flexible, general-purpose language understanding (Gain et al., 2025). LLMs like GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2022), and GPT-4 demonstrate strong multilingual capabilities without explicit training on translation tasks. LLM-based translation leverages in-context learning, where the model is prompted with examples or instructions to perform translation on the fly. This approach has shown competitive performance in zero-shot and few-shot learning scenarios (Vilar et al., 2023; Luo et al., 2025), especially for high-resource language pairs. Unlike traditional neural machine translation (NMT), which requires retraining or fine-tuning for each new domain or language, LLMs can generalize across tasks and languages with minimal additional data.
@@
More recently, the emergence of LLMs has opened new possibilities for Multi-Agent Systems (MAS). In LLM-based multi-agent systems, each agent is instantiated as an LLM-powered entity capable of natural language reasoning, planning, and collaboration. Systems such as AutoGPT (Yang et al., 2023), CAMEL (Li et al., 2023), and AutoGen (Dibia et al., 2024) demonstrate that LLM agents can simulate diverse roles and complete complex tasks through dialogue-based coordination. A growing number of studies explore the use of multi-agent systems for translation-related tasks. Notably, MAS has emerged as a promising solution for document-level translation (Wang et al., 2024), a long-standing challenge in MT.
@@
Formatted text translation involves translating documents that contain structural or semantic markup, such as LaTeX and XML. These formats often interleave natural language with commands, tags, or tokens that encode formatting, layout, or semantic annotations. Although some recent efforts have been made in this direction Kleidermacher and Zou (2025); Khan (2025), formatted text translation still faces two major challenges. The first is the lack of a robust, general-purpose system specifically designed for translating formatted content. Currently, only a few proprietary tools, such as Youdao and Baidu, offer relatively effective solutions. While open-source tools like MathTranslate***https://github.com/SUSYUSTC/MathTranslate and GPT-Academic†††https://github.com/binary-husky/gpt_academic have received positive feedback, they still lag behind commercial systems in overall performance. The second is the lack of a sound, formatted text translation evaluation technique. As traditional BLEU or COMET scores do not cover format correctness or tag retention. Therefore, it is imperative to develop a new evaluation technique for structure-aware translation.
@@ 3 System Design
The key architecture of LaTeXTrans is a multi-agent coordination designed for translating structured LaTeX documents. It consists of three modules: the Parser, the Translation Module, and the Generation Module. The design and functionality of each component are described in detail below.
@@ 3.3 Generation Module
The generation module is responsible for reassembling the translation units into structured LaTeX documents and compiling the structured LaTeX documents into PDF files using specific compilers (e.g. pdfLaTeX and XeLaTeX).
@@ 4 Experiment
| System | En-Zh | En-Ja | ||||||
| Cometkiwi () | LLM-score () | FC-score () | Cost () | Cometkiwi () | LLM-score () | FC-score () | Cost () | |
| NiuTrans | 64.69 | 7.93 | 60.72 | - | 65.49 | 8.19 | 27.48 | - |
| Google Translate | 46.23 | 5.93 | 51.00 | - | 56.21 | 7.01 | 50.00 | - |
| LLaMA-3.1-8b | 42.89 | 2.92 | 49.40 | - | 44.49 | 3.32 | 60.92 | - |
| Qwen-3-8b | 45.55 | 7.87 | 48.68 | - | 46.20 | 6.80 | 49.52 | - |
| Qwen-3-14b | 68.18 | 8.76 | 65.63 | - | 72.84 | 8.66 | 61.88 | - |
| DeepSeek-V3 | 67.26 | 9.02 | 63.68 | $0.02 | 72.17 | 9.00 | 63.96 | $0.03 |
| GPT-4o | 67.22 | 8.58 | 58.32 | $0.13 | 71.16 | 8.91 | 56.92 | $0.11 |
| LaTeXTransQwen-3-14b | 71.37 | 8.97 | 71.20 | - | 74.68 | 8.51 | 59.84 | - |
| LaTeXTransDeepSeek-V3 | 73.48 | 9.01 | 70.52 | $0.10 | 75.39 | 8.89 | 66.52 | $0.13 |
| LaTeXTransGPT-4o | 73.59 | 8.92 | 71.52 | $0.35 | 74.47 | 8.93 | 64.92 | $0.45 |
@@ 4.1 Settings
Since no publicly available LaTeX document dataset currently exists, we constructed our test set by selecting the TeX sources of 50 English academic papers from the arXiv repository. The chosen papers include both long and short articles, many of which contain complex formulas and figures, ensuring structural diversity and complexity in the LaTeX content. Further experimental details are provided in Appendix A.
@@
Our baselines are categorized into two groups: traditional MT systems and LLM-based translation systems. For the former, we selected NiuTrans and Google Translate as representative systems. For the latter, we evaluated five strong LLMs, including both open-source and proprietary models: LLaMA-3.1-8B Grattafiori et al. (2024), Qwen-3-8B Yang et al. (2025), Qwen-3-14B, DeepSeek-V3 Liu et al. (2024), and GPT-4o Hurst et al. (2024). Among these, Qwen-3-14B, DeepSeek-V3, and GPT-4o were further used as the backbone models for agents in LaTeXTrans.
4.2 Evaluation Metrics
We conducted a comprehensive assessment of our system from two dimensions: translation quality and format retention ability.
@@ 4.2 Evaluation Metrics
Because high-quality reference translations for LaTeX documents require expert-level annotation, we adopted wmt22-cometkiwi-da Rei et al. (2022), a reference-free evaluation metric (denoted as Cometkiwi), to assess the translation quality of LaTeX documents. Furthermore, we employed GPT-4o as an automatic evaluator to further assess translation quality across multiple dimensions, guided by carefully designed system prompts. The evaluation covered four aspects: Faithfulness, Fluency, Terminology Consistency, and Coherence, where each was rated on a scale from 0 to 10. An overall score was then synthesized by GPT-4o based on the individual scores across these dimensions (denoted as LLM-score).
Format Retention Ability.
Whether the labels are completely retained is an important manifestation of the ability of the formatted text translation system. However, at present, there is no universal indicator to evaluate the format retention ability of models or systems during the translation process. Therefore, for LaTeX documents, we have designed a new evaluation metric, Format Consistency Score (denoted as FC-score), to assess the retention ability of our system for LaTeX labels during the translation process. We can compute the FC-score by
where is the initial score before the rewards and penalties, is the penalty coefficient per error, is the penalty coefficient per warning, is the reward for successful compilation. and are numbers of errors and warnings, indicates whether the LaTeX document compiled successfully. We then clip the score to the valid range , and are the upper bound and lower bound of the score (e.g. 0~100).
4.3 Results
We evaluate our LaTeXTrans system on two translation tasks: English-to-Chinese (En-Zh) and English-to-Japanese (En-Ja). The results, shown in Table 1, demonstrate that LaTeXTrans consistently outperforms both the NMT and Single-Agent baselines across all evaluation metrics, including COMETkiwi and FC-score. In terms of translation quality, LaTeXTrans demonstrates substantial improvements in FC-score (71.52 vs. 58.32 for the En-Zh task and 70.52 vs. 63.68 for the En-Ja task), indicating significantly better preservation of LaTeX formatting during translation. Moreover, when powered by GPT-4o as the backbone model, LaTeXTrans achieves the highest scores across all three evaluation metrics—COMETkiwi, LLM-score, and FC-score—underscoring its strong overall translation performance on structured LaTeX documents. In terms of translation cost, LaTeXTrans delivers superior performance without incurring a substantial increase in computational expense compared to other LLM-based translation systems, making it well-suited for large-scale deployment in real-world applications.
@@ 4.4 Ablation Study
Table 2 presents an ablation study on the En–Zh task using GPT-4o and DeepSeek-V3 as backbone models. Introducing the Parser module significantly improves both COMETkiwi and FC-score, indicating that the placeholder substitution strategy enhances translation quality and label preservation. Adding the Validator module further boosts overall performance, although a slight drop in LLM-score is observed with DeepSeek-V3. We hypothesize that this is due to the Validator enforcing strict tag retention through iterative checks, which may restrict the Translator and slightly impact fluency. Finally, incorporating the Summarizer and Terminology Extractor improves the LLM-score, reflecting better cross-paragraph coherence. However, slight declines in COMETkiwi and FC-score suggest that these improvements may not be fully captured by COMETkiwi. A detailed analysis with a case study is provided in Section 4.4.1.
| Setting | GPT-4o | DeepSeek-V3 | ||||
| Cometkiwi | LLM-score | FC-score | Cometkiwi | LLM-score | FC-score | |
| SA. (Baseline) | 67.22 | 8.58 | 58.32 | 67.26 | 9.02 | 63.68 |
| SA. + P. | 74.47 | 8.89 | 69.64 | 74.39 | 9.03 | 70.08 |
| SA. + P. + V. | 74.57 | 8.91 | 71.76 | 74.42 | 8.94 | 70.80 |
| SA. + P. + V. + S. | 74.06 | 8.95 | 71.64 | 74.02 | 9.05 | 70.68 |
| SA. + P. + V. + S. + TE. | 73.59 | 8.93 | 71.52 | 73.48 | 9.01 | 70.52 |
4.4.1 Translation consistency
We present a case study of the En-Zh task from our test set to demonstrate that our system does indeed perform better in terms of translation consistency, as shown in Figure 3. In this case, the terminology translation of LaTeXTrans remains consistent across the three sections. In contrast, the baseline method finds it difficult to maintain such consistency. This indicates that our system can maintain excellent consistency throughout the entire translation process.

### v3

@@ Abstract
Despite the remarkable progress of modern machine translation (MT) systems on general-domain texts, translating structured LaTeX-formatted documents remains a significant challenge. These documents typically interleave natural language with domain-specific syntax, such as mathematical equations, tables, figures, and cross-references, all of which must be accurately preserved to maintain semantic integrity and compilability. In this paper, we introduce LaTeXTrans, a collaborative multi-agent system designed to address this challenge. LaTeXTrans ensures format preservation, structural fidelity, and terminology consistency through six specialized agents: 1) a Parser that decomposes LaTeX into translation-friendly units via placeholder substitution and syntax filtering; 2) a Translator, Validator, Summarizer, and Terminology Extractor that work collaboratively to ensure context-aware, self-correcting, and terminology-consistent translations; 3) a Generator that reconstructs the translated content into well-structured LaTeX documents. Experimental results show that LaTeXTrans outperforms mainstream MT systems in both translation accuracy and structural preservation. The source code***https://github.com/NiuTrans/LaTeXTrans, the online demonstration platform†††https://latextrans.online, and a demo video‡‡‡https://youtu.be/-ODRUTE-VU8 are publicly available.
@@
Ziming Zhu1††thanks: Authors contributed equally., Chenglong Wang1∗, Haosong Xu1, Shunjie Xing1, Yifu Huo1, Fengning Tian2, Quan Du2, Di Yang1,2, Chunliang Zhang1,2, Tong Xiao1,2††thanks: Corresponding author. and Jingbo Zhu1,2 1School of Computer Science and Engineering, Northeastern University, Shenyang, China 2NiuTrans Research, Shenyang, China {zhuzm0721, clwang1119}@gmail.com, {xiaotong, zhujingbo}@mail.neu.edu.cn
@@ 1 Introduction
LaTeX Urban (1986) is a widely used macro system built on TeX for typesetting complex and structured documents, and has become the standard infrastructure for scholarly publishing worldwide. Although LaTeX itself is language-agnostic, academic communication remains overwhelmingly dominated by English: nearly 98% of scientific papers are published in English Kleidermacher and Zou (2025), while only about 5% of the global population is proficient in the language as of 2025 Dwyer (2025). This extreme asymmetry imposes a significant burden on most researchers. Non-native English speakers must not only master domain knowledge but also navigate linguistically demanding and structurally complex LaTeX documents in a foreign language.
To alleviate this burden, the most straightforward approach is to first compile LaTeX source files into a PDF and then apply machine translation directly at the PDF level (a.k.a. PDF translation). However, this approach often results in incomplete formatting due to errors in PDF parsing. An alternative line of research advocates translating directly at the LaTeX source level before recompilation into the target language. Although prior work has demonstrated improved format preservation compared to PDF translation Hoy et al. (2025), such systems are typically limited to short and well-structured LaTeX source inputs. In practice, however, real-world LaTeX projects are rarely monolithic or syntactically simple. They often consist of multiple interdependent source files, user-defined commands, and domain-specific macros.
Given these limitations, there remains a strong demand for a translation system capable of handling real-world LaTeX projects in their full structural complexity. To this end, we propose LaTeXTrans, a collaborative multi-agent framework that directly translates raw LaTeX source files while preserving structural fidelity and semantic integrity. Unlike prior approaches that focus on isolated snippets, LaTeXTrans is designed to robustly handle multi-file architectures, user-defined commands, and macro-intensive environments. It consists of three modules and six specialized agents:
Parsing Module: Responsible for fine-grained analysis of LaTeX-formatted documents. To handle the structural complexity of LaTeX projects, we design a Parser agent equipped with a placeholder mechanism and a syntax filter, which together decompose the source into manageable translation units.
@@
To evaluate LaTeXTrans, we construct a multi-domain, multilingual benchmark for LaTeX source-level translation using real-world TeX projects from arXiv. We compare our system with representative baselines for formatted-text translation, assessing both structural preservation and translation accuracy. The results demonstrate that LaTeXTrans robustly handles complex LaTeX project architectures, maintains syntactic and structural consistency during translation, and consistently generates successfully compiled target-language PDFs. In the En–Zh setting, LaTeXTrans achieves a one-pass compilation success rate of 97% across all test samples, with an average formatting error rate below 0.5 per document.
@@ 2 Related works
The emergence of large language models (LLMs) has introduced a new paradigm for machine translation, moving beyond traditional supervised training on parallel corpora toward more general-purpose language understanding Gain et al. (2025). Models such as GPT-3 Brown et al. (2020), PaLM Chowdhery et al. (2022), and GPT-4 exhibit strong multilingual capabilities without task-specific translation training. By leveraging in-context learning, LLM-based translation performs competitively in zero-shot and few-shot settings (Vilar et al., 2023; Luo et al., 2025), particularly for high-resource language pairs. Unlike traditional neural machine translation (NMT), which often requires domain- or language-specific retraining, LLMs demonstrate broader cross-task and cross-lingual generalization with minimal additional data.
@@
More recently, the emergence of LLMs has opened new possibilities for multi-agent systems (MAS). In LLM-based MAS, each agent is instantiated as an LLM-powered entity capable of natural language reasoning, planning, and collaboration. Systems such as AutoGPT (Yang et al., 2023), CAMEL (Li et al., 2023), and AutoGen (Dibia et al., 2024) demonstrate that LLM agents can simulate diverse roles and complete complex tasks through dialogue-based coordination. A growing number of studies explore the use of MAS for translation-related tasks. Notably, MAS has emerged as a promising solution for document-level translation (Wang et al., 2024), a long-standing challenge in MT.
@@
Formatted text translation concerns documents that contain structural or semantic markup, such as LaTeX and XML, where natural language content is interleaved with commands, tags, and tokens encoding layout or semantic information Kleidermacher and Zou (2025); Khan (2025). Research in this field faces two primary challenges. Firstly, there is a lack of robust, general-purpose systems specifically designed for translating formatted documents. While several proprietary tools (e.g., Youdao and Baidu) provide practical solutions, open-source systems such as MathTranslate§§§https://github.com/SUSYUSTC/MathTranslate and GPT-Academic¶¶¶https://github.com/binary-husky/gpt_academic offer limited performance and scalability. Secondly, evaluation methodologies for formatted text translation remain inadequate. Conventional metrics such as BLEU Papineni et al. (2002) and COMET Rei et al. (2020) focus on linguistic quality but fail to assess structural correctness and compilability.
@@ 3 System Design
The main architecture of LaTeXTrans is a multi-agent coordination designed for translating LaTeX source files. It consists of three modules: the Parser, the Translation Module, and the Generation Module. The design and functionality of each component are described in detail below.
@@ 3.3 Generation Module
The generation module automatically selects the appropriate compilation engine (e.g., pdfLaTeX or XeLaTeX) based on the document characteristics, and subsequently compiles the structured and translated LaTeX source files into PDF outputs.
4 System Deployment and Usage
4.1 Command-Line Tool
LaTeXTrans supports local deployment through a command-line interface. After configuring the API credentials, we can run fully automated end-to-end translation. By providing an arXiv paper ID, the system retrieves the associated LaTeX project and produces a translated PDF in a single command:
LaTeXTrans --arxiv xxxx.xxxxx
Beyond API setup, LaTeXTrans supports flexible customization, including target language selection and the integration of user-defined terminology dictionaries. Upon completion, the system generates both the translated PDF and the complete translated LaTeX source project. Additionally, detailed translation and compilation logs are provided to facilitate debugging and ensure transparency.
4.2 Online Platform
To further enhance accessibility, we provide a web-based platform built upon LaTeXTrans for LaTeX document translation. The platform supports three input modes: (1) automatic retrieval via arXiv paper IDs, (2) batch processing of multiple IDs, and (3) uploading local LaTeX projects. During execution, users can monitor real-time translation and compilation logs. Additionally, the platform offers bilingual PDF comparison to facilitate side-by-side reading. The usage workflow of the online platform is illustrated in Figure 3.
@@ 5.1 Settings
As no publicly available benchmark currently exists for LaTeX source-level document translation, we construct a test set comprising the LaTeX source files of 100 English arXiv papers∥∥∥Although the test set contains 100 papers, this scale is sufficient for LaTeX-based evaluation, as each paper typically spans multiple pages and contains rich structural elements, resulting in substantial document-level complexity.. The dataset includes 50 papers from computer science, 30 from physics, and 20 from mathematics. To further stress-test the system, we deliberately include 10 long survey papers characterized by complex layouts and extended document structures. This design ensures sufficient scale, domain diversity, and structural complexity, enabling a comprehensive and realistic evaluation of our approach.
@@
To comprehensively evaluate both translation quality and format preservation, we categorize the baselines into two groups: section-level direct translation and formatted-text translation systems. For the former, we selected NiuTrans and Google Translate as traditional MT systems, and Qwen-3-14B Yang et al. (2025), DeepSeek-V3 Liu et al. (2024), and GPT-4o Hurst et al. (2024) as LLM-based translation systems. For the latter, we selected gpt-academic as the baseline. When evaluating section-level direct translation systems, we segment LaTeX documents into section-level units and translate them directly without structural parsing or iterative validation.
| Subject | En-Zh | En-Ja | En-Ko | |||
| CSR () | FEC () | CSR () | FEC () | CSR () | FEC () | |
| Physics | 100% | 0.37 | 93% | 0.43 | 90% | 0.47 |
| Mathematics | 100% | 1.15 | 95% | 1.30 | 100% | 1.35 |
| Computer Science | 94% | 0.18 | 76% | 0.20 | 86% | 0.25 |
| Total | 97% | 0.43 | 85% | 0.49 | 90% | 0.54 |
5.2 Main Results
Structural Preservation.
To evaluate the structural performance of LaTeXTrans, we report the Compilation Success Rate (CSR) and the Formatting Error Count (FEC). CSR measures the percentage of translated LaTeX projects that compile successfully without manual intervention, while FEC represents the average number of formatting errors per document. Detailed error categories and representative examples are provided in the Appendix B. We evaluate LaTeXTrans on three translation tasks: English-Chinese (En–Zh), English–Japanese (En–Ja), and English–Korean (En–Ko). As shown in Table 1, we find that LaTeXTrans can achieve a CSR of 97% in the En–Zh setting and consistently high CSR scores in the En–Ja and En–Ko tasks, demonstrating strong structural reliability across diverse LaTeX projects. Notably, CSR reaches 100% in the physics and mathematics domains. Beyond compilation success, LaTeXTrans can maintain consistently low FEC values across all tasks. Note that the slightly higher FEC observed in mathematics is primarily due to the dense and complex mathematical environments, which increase sensitivity to formatting inconsistencies during reconstruction.
| System | En-Zh | En-Ja | ||
| Cometkiwi | LLM-score | Cometkiwi | LLM-score | |
| NiuTrans | 64.69 | 7.93 | 65.49 | 8.19 |
| Google Translate | 46.23 | 5.93 | 56.21 | 7.01 |
| Qwen-3-14b | 68.18 | 8.76 | 72.84 | 8.66 |
| LaTeXTrans Qwen-3-14b | 71.37 | 8.97 | 74.68 | 8.51 |
| DeepSeek-V3 | 67.26 | 9.02 | 72.17 | 9.00 |
| LaTeXTrans DeepSeek-V3 | 73.48 | 9.01 | 75.39 | 8.89 |
| GPT-4o | 67.22 | 8.58 | 71.16 | 8.91 |
| LaTeXTrans GPT-4o | 73.59 | 8.92 | 74.47 | 8.93 |
| System | Total (70) | Mathematics (20) | Computer Science (50) | ||||
| A | B | C | A | B | C | ||
| LaTeXTrans | 67 | 14 | 3 | 3 | 45 | 1 | 1 |
| gpt-academic | 69 | 9 | 6 | 5 | 40 | 7 | 2 |
@@ 5.2 Main Results
We evaluate LaTeXTrans using COMETkiwi Rei et al. (2022) and LLM-score. Detailed evaluation settings are provided in Appendix A. The results are listed in Table 2. First, compared with traditional MT systems, LaTeXTrans demonstrates clear performance advantages, highlighting the effectiveness of our LaTeX-specific optimization strategies. When compared to LLM-based translation approaches, our system achieves comparable performance while preserving structural fidelity, and even surpasses them on certain metrics. We attribute these gains to our fine-grained parsing strategy and placeholder-based isolation of non-textual elements.
Human Evaluation.
To complement automatic structural metrics, we conduct a human evaluation on the En–Zh translation task to assess format preservation quality. Translated documents are categorized into three levels based on formatting quality (see Appendix A for details): (A) perfect preservation of the original format, (B) minor formatting errors that do not affect readability, and (C) severe errors that significantly impair reading. As shown in Table 3, LaTeXTrans achieves a substantially higher proportion of level-A outputs compared to GPT-Academic, demonstrating superior format preservation. These findings indicate that LaTeXTrans not only improves automatic structural metrics but also delivers visibly higher formatting quality from a human evaluation.
@@ 5.3 Ablation Study
Table 4 presents an ablation study on the En–Zh task. Introducing the Parser module significantly improves COMETkiwi, indicating that the placeholder substitution strategy enhances translation quality. Adding the Validator module further boosts overall performance, although a slight drop in LLM-score is observed with DeepSeek-V3. We hypothesize that this is due to the Validator enforcing strict tag retention through iterative checks, which may restrict the Translator and slightly impact fluency. Finally, incorporating the Summarizer and Terminology Extractor improves the LLM-score, reflecting better cross-paragraph coherence.
| Setting | GPT-4o | DeepSeek-V3 | ||
| Cometkiwi | LLM-score | Cometkiwi | LLM-score | |
| SA. (Baseline) | 67.22 | 8.58 | 67.26 | 9.02 |
| SA. + P. | 74.47 | 8.89 | 74.39 | 9.03 |
| SA. + P. + V. | 74.57 | 8.91 | 74.42 | 8.94 |
| SA. + P. + V. + S. | 74.06 | 8.95 | 74.02 | 9.05 |
| SA. + P. + V. + S. + TE. | 73.59 | 8.93 | 73.48 | 9.01 |